Andrew and I have disassembled our electronics rack.
It will rise again in three days, in fulfillment of the scriptures.
Gone is the messy wire rail with Phoenix Contact fuses. Gone is the VME circuit card holder. Gone is the desktop computer acromag1 (10.0.1.33). Gone is the oscilloscope 360noscope (10.0.1.71). Gone are the two HP variable power supplies.
Still there are the KEPCO power supplies, the quiet APC laser power supply, the crappy acromag card panel (soon to be replaced thnx 2 awade), the two TTFSS panels (D0902048), the 36.0 and 37.0 MHz crystal oscillators panel (D1600008), and a nice BNC breakout panel.
To be added are new Sorenson power supplies, nice +/- 24 V and +/- 18 V LIGO DC power rails, and more supports custom made by awade as I write this.
While removing the messy wire rail, I noticed some broken connections. In the RackResurrection.pdf attached, the broken connections are shown in pictures 9 and 12. Picture 9 shows the + 24 V power supply with two ground wires detached. Picture 12 shows a wire from the + 17 V rail to one of the TTFSS panels detached from the wire rail. Unclear if these broken connections are causing our FSS box issues, it surely can't help. The broken connection in Picture 12 could have happened during disassembly, but the + 24 V power supply one was there when I started.
The LIGO DC power rail shovel connectors are not compatible with the KEPCO power supplies, they are too big. I have requested assistance from Gautam in locating the correct shovels.
In addition to the rack's demise, the ceiling cable ladder has been purged. Any long cable that can be rerouted has been rerouted. I added a plastic bridge for wiring BCNs to the cable ladder from the middle of the optics bench.
The power surge protectors on the underside of the table supports have been completely unplugged, moved, and reattached. The rails on the north side have been daisy chained together, and then plugged into the APC power supply. The same was down for the south. The north surge protectors have the lasers plugged into them, so quiet, low intensity fluctuation power is necessary there.
The desktop computer acromag1 has been placed underneath the clean bench for the time being. We have acquired an extra monitor, so soon we will have acromag1, ws3, and awade's raspberry pi running in our lab (LAN party).
The laser mains supply box has been removed from the cable ladder and placed underneath the optics bench. Safety expert awade has arranged things such that the emergency laser stop button and laser safety lights will be properly connected and actually work to keep us safe.
The PLL equipment is sitting on the rolling shelf. We want to put the Marconi and preamp on top of the rack eventually.
We also must redo the crappy photodiode connectors and crappy temperature sensor wire mess. The temperature sensor box must replace this whole clipdoodle power supply nonsense we had.
The first thing we did this afternoon was take down all the wrapping and turn on the HEPA filters, the clean bench, and the ion pump. The ion pump power supply has been placed where the PLL electronics used to be, above the transmission PDs. We removed all electronics from under the bench, cleaned the floor with awades robot, and cleaned the top of the rack from dust.
Finishing the rack electronics will be a week of effort, but at the end of the week we want to have everything relocked with a beatnote, with vac can thermal controls on the rack and working.